OpenStudy (anonymous):

Lana is the oldest of four sisters. Her youngest sister is half her age. The other two sisters are twins 2 years younger than Lana. When the four sisters' ages are added together, the result is 45. If the youngest sister is n years old, what is n? will give meadel

OpenStudy (anonymous):

I will give a medal

OpenStudy (anonymous):

If the youngest sister is \(n\), then the oldest is \(2n\) because the youngest sister is half her age.

OpenStudy (kash_thesmartguy):

http://52.0.184.118/updates/53f75c68e4b0ae4669f46685 -link might help you!

OpenStudy (anonymous):

Then two sisters are twins and they are 2 years younger than Lana.

OpenStudy (anonymous):

We know Lana is \(2n\), so the two sisters are both \(2n-2\).

OpenStudy (anonymous):

Set up the equation. Lana+ twin +twin+youngest =45

OpenStudy (anonymous):

so what age is the youngest?

OpenStudy (anonymous):

The youngest is n. I put that first. Read it.

OpenStudy (anonymous):

so what is n?

OpenStudy (anonymous):

\[2n+2n-2+2n-2+n=45\]Now solve for n.

OpenStudy (anonymous):

I don't give out answers. YOU work for it.

OpenStudy (anonymous):

so is it 7n-4?

OpenStudy (anonymous):

Yes, but you need to find for n, right?

OpenStudy (anonymous):

so 7n-4=5

OpenStudy (anonymous):

\[\huge 7n-4=45\]

OpenStudy (anonymous):

Add 4 on both sides. Then divide 7 on both sides.

OpenStudy (anonymous):

so it is 7?

OpenStudy (anonymous):

yes